Stability of solving systems Differential Equations of special form in R2

Nasiba Raupova Samarkand State University

Scrutinized common-line system of differential equations with definition At work we are investigating the stability of solving dimensional system right parts that are common-line functions.There is a function of Lyapunova and exposed enough conditions for the given function.